A 1.00-L gas sample at 100.°C and 600. torr contains 49.0% helium and 51.0% xenon by mass. What are the partial pressures of the individual gases?
1
Expert's answer
2017-12-01T02:00:07-0500
100 g of gas = 49 g of He and 51 g Xe
n(He) = 49 / 4 = 12.25 mol
n(Xe) = 51 / 131.29 = 0.388 mol
n(He) + n(Xe) = 12.25 + 0.388 = 12.638 mol
pHe / 600 = 12.25 / 12.638
pHe = 581.5 torr
pXe / 600 = 0.388 / 12.638
pXe = 18.4 torr